OpenStudy (rainbow_rocks03):

PLZ HELP ME I will medal+fan you for BEST response. Solve for the variable.

OpenStudy (rainbow_rocks03):



OpenStudy (rainbow_rocks03):



OpenStudy (anonymous):

You solve this problem by cross multiplying and dividing, do you know how to do that?

OpenStudy (rainbow_rocks03):

6*3=18

OpenStudy (rainbow_rocks03):

Is the answer C.?

OpenStudy (anonymous):

thats half of it, you also have to multiply the G and the 36 to get: G*36=6*3. Then you divide the left and ride sides by 36 to be left with your answer for G. Try that, no C is not correct

OpenStudy (rainbow_rocks03):

Like this? :6*3=18/36=0.5?

OpenStudy (anonymous):

Yep that's exactly right, you just solved for G congratulations!

OpenStudy (rainbow_rocks03):

so the answer is B.

OpenStudy (rainbow_rocks03):

YAY!

OpenStudy (anonymous):

Correct good job!